# Superfluid Dark Matter

### v1

@@ 1 Introduction
Cosmological and astrophysical data demonstrate that Cold Dark Matter (CDM) accounts for roughly a quarter of the total energy density of the Universe [1]. Within the $\Lambda$CDM framework, dark matter (DM) is modeled as a pressureless cosmological fluid, leading to remarkable agreement with large-scale structure observations [2] and the temperature anisotropies of the cosmic microwave background [1]. Despite its success on cosmological scales, the microscopic nature of DM remains elusive. Moreover, several persistent anomalies in galactic dynamics lack a natural—or at least well-understood—explanation within the standard paradigm, prompting interest in models with richer microscopic properties that could impact astrophysical behavior. For comprehensive reviews, see Refs. [3, 4, 5].
@@ 1.1 Cold dark matter and small scale challenges
In the standard $\Lambda$CDM model, the collapse of the primordial fluctuations leads to the formation of virialized halos. These describe overdense DM regions with enclosed mass
@@
However, numerical simulations have encountered challenges in reproducing astrophysical observations on galactic scales. While many of these small-scale issues are complicated by the uncertainties inherent in modeling baryonic physics and observational systematics, certain discrepancies have emerged over the past years as significant and persistent puzzles for the $\Lambda$CDM paradigm.
@@
The overprediction of the number of halos, commonly known as the missing satellites problem, has long represented a significant challenge for $\Lambda$CDM. However, this issue has largely been alleviated in recent years. In hierarchical structure formation models, DM halos are expected to host a population of smaller subhalos, with the halo mass function quantifying their abundance within specific mass intervals. This yields a robust prediction within the $\Lambda$CDM framework: as the mass of subhalos increases, their number decreases monotonically across all observable mass ranges. In particular, simulations predict that Milky Way-like galaxies should host approximately a thousand subhalos with masses exceeding $10^{7}M_{\odot}$ [41, 42]—masses large enough to potentially host visible galaxies. Initially this prediction appeared to be in stark contrast with the relatively small number of satellites observed around the Milky Way. However, beginning in the mid-2000s, improved observational campaigns have led to the discovery of numerous faint satellite galaxies in both the Milky Way and Andromeda, resolving halos with masses as small as approximately $300\,M_{\odot}$ [43, 44]. The number of detected satellites is expected to continue increasing with future surveys.
@@
These issues constitute some of the most prominent small-scale challenges to the $\Lambda$CDM model. For a more detailed discussion, we refer the interested reader to [5].
@@ 1.2 Beyond the cold dark matter paradigm
One of the simplest modifications to the DM paradigm that addresses small-scale discrepancies is warm DM [71]. In this model, DM particles have a small mass, resulting in a thermal velocity dispersion that alters their behavior on galactic scales while preserving the successful large-scale predictions of the CDM model. Even with modest velocities, DM particles can free-stream out of potential wells, suppressing the formation of small-scale structures and mitigating several associated problems [72, 73, 74, 75, 76].
Another model motivated by these discrepancies is self-interacting DM [77, 78], in which DM particles interact with one another with a cross section of order $\sigma/m\lesssim{\rm cm}^{2}/{\rm g}$. Self-interactions redistribute DM from the inner to the outer regions of halos, softening the central cusp of the NFW profile and suppressing small-scale structure formation [79, 80]. This model can reproduce the observed diversity in rotation curve shapes of dwarf galaxies [81, 82, 83], and alleviates the too-big-to-fail problem.
@@
Yet another scenario that has gained significant attention in recent years is fuzzy DM, due to its rich phenomenology on small scales [84, 85]. In this model, DM consists of ultralight bosons with masses lighter than $10^{-20}\,{\rm eV}$. The extremely small mass invalidates a classical fluid description and allows DM to form a condensate on galactic scales through the balance between quantum pressure and gravity. This wave-like behavior gives rise to several distinctive phenomena, including the suppression of substructures below the de Broglie wavelength and the formation of cores in the inner regions of halos, with characteristic sizes set by quantum pressure and gravity [86, 87, 88]. On cosmological scales, however, fuzzy DM behaves similarly to CDM, preserving the large-scale successes of the standard model. The fuzzy DM model is constrained by Lyman-$\alpha$ forest observations [89, 90, 91, 92], which place a lower bound of $m>2\times 10^{-20}$ eV at 95% confidence level, as well as dwarf galaxies [93, 94, 95] and rotation curve data [96]. For recent reviews on fuzzy DM, see Refs. [97, 98, 4, 99].
@@
The extremely light mass range considered in fuzzy DM models requires that gravitational collapse be counterbalanced by quantum pressure. However, one can envision alternatives involving heavier DM candidates, extending up to the eV mass range. A prime example is the superfluid DM model, which postulates sub-eV DM particles with self-interactions capable of undergoing a superfluid phase transition in galactic halos. [^1] The formation of superfluids in galaxies can also be achieved through fermions with attractive interactions via Bardeen-Cooper-Schrieffer pairing. This can relax the Tremaine-Gunn lower bound [100] on the DM fermion mass. For developments in this direction, see Refs. [101, 102, 103]. This framework allows for the formation of coherent superfluid cores, stabilized by the balance between self-interactions and gravity [6, 7, 8]. It offers a distinct approach to addressing the small-scale challenges mentioned earlier and admits, in principle, microscopic realizations. [^2] Another ultralight DM model, named self-interacting fuzzy DM, shares the same mass range as standard fuzzy DM but incorporates weak self-interactions [104, 105, 106, 107]. The main difference with SDM lies in the coherence length of the system. Although both frameworks involve self-interacting ultralight particles, self-interactions in the fuzzy case are not efficient enough to bring the system to equilibrium. As a result, the coherence length remains approximately set by quantum pressure, as in the non-interacting fuzzy DM scenario, in contrast to SDM, where self-interactions govern the macroscopic coherence properties. For earlier reviews on the SDM scenario, see Refs. [97, 4]. A schematic summary of the DM models discussed is shown in Fig. 1.
@@
In this review, we focus on the SDM scenario, summarizing its distinctive phenomenology on galactic scales and its potential to address small-scale challenges. Section 2 introduces the phenomenon of superfluidity, outlining its theoretical foundations within the frameworks of effective field theory and hydrodynamics. We also discuss finite-temperature effects, the interaction of the superfluid with external perturbations, and the role of gravity. Section 3 focuses on DM superfluidity in galactic environments, with particular attention to the formation of superfluid halos and their observable phenomenology. Section 4 explores its cosmological aspects, including possible production mechanisms and the subsequent evolution of the condensate. Section 5 provides an overview of the phenomenological implications of the SDM framework in both astrophysical and cosmological contexts, with particular emphasis on vortex formation within halos, the development of DM spikes around BHs, galaxy rotation curves, and the dynamics of merging galaxy clusters. Section 6 examines the phenomenon of dynamical friction and its suppression in the presence of a superfluid, in contrast to the expectations from standard CDM. We present the theoretical foundations of this effect using both a quasiparticle description and a hydrodynamical approach. An application to the evolution of BH binaries is also discussed. Section 7 explores the possibility of long-range forces arising between baryons immersed in the DM superfluid, by coupling them to the constituent degrees of freedom. We present two illustrative examples capable of generating such interactions. Finally, Section 8 summarizes the main results of the review.
@@ 2 Superfluidity and its effective description
where $c_{s}$ is the sound speed of the superfluid, expressed in terms of its pressure $P$ and density $\rho$ [^3] As usual, $\omega_{k}$ and $k$ denote the energy and momentum of the wave, respectively.
@@
where $E_{\text{\tiny kin}}$ is the kinetic energy of the superfluid. For the superfluid to be slowed down by the passage of the perturber, a certain amount of kinetic energy has to be damped by excited phonons. This is possible if the sum of the second and third terms of Eq. (2.2) is negative. By plugging in the phonon spectrum in Eq. (2.1), we have dissipation if $c_{s}k+\vec{k}\cdot\vec{v}_{s}\,<0$, which requires
@@ 2.1 Simplest example: superfluidity in a quartic theory
Here, $v$ describes the vacuum expectation value of the field, while $\mu$ is the relativistic chemical potential of the system, which is chosen in such a way that the field configuration (2.5) satisfies the equation of motion. In the non-relativistic limit, we can expand $\mu$ to obtain
@@
Several approximations can be considered to simplify the analysis. First, since we aim to study non-relativistic superfluids, we can drop the $\dot{h}^{2}$ and $\dot{\pi}^{2}$ terms. This implies that the typical energy scales of interest are smaller than the mass $m$ of the scalar. Second, we can consider a leading-order approximation in spatial derivatives of $h$ by dropping the gradient term $\big{(}\vec{\nabla}h\big{)}^{2}$. This is possible if we are interested in energy scales smaller than $\mu_{\text{\tiny NR}}$. In the context of DM superfluidity, the shortcomings of this approximation were discussed for dissipative processes [113] and in the analysis of the gravitational stability of the superfluid phase [114].
@@
Lastly, we comment on the requirement of repulsive self-interactions for superfluidity. If the theory had negative quartic coupling, $\lambda_{4}<0$, the long-wavelength fluctuations of the homogeneous solution would have an imaginary sound speed and thus would be unstable. An example of this class of theories is given by axions, and it implies that such particles fail to generate long-range correlations due to their attractive potential [115]. The challenges associated with developing a theory of pseudo Nambu-Goldstone bosons with repulsive self-interactions have also been mentioned in [116].
@@ 2.4 Gross–Pitaevskii equation and the hydrodynamical formulation
Substituting back into (2.30), and ignoring terms quadratic in time derivatives, we obtain the Schrödinger-like equation,

### v2

@@ 1 Introduction
[…]
@@ 1.1 Cold dark matter and small scale challenges
[…]
@@
[…]
@@
[…]
@@
[…]
@@ 1.2 Beyond the cold dark matter paradigm
One of the simplest modifications to the DM paradigm that addresses small-scale discrepancies is warm DM [130]. In this model, DM particles possess a small but non-negligible mass, typically in the range of $m\sim\mathrm{keV}$, leading to a residual thermal velocity dispersion that suppresses the growth of density perturbations below a characteristic free-streaming scale. As a result, warm DM behaves similarly to CDM on large cosmological scales while exhibiting distinct dynamics on subgalactic scales. Even with modest velocities, DM particles can erase primordial fluctuations by free-streaming out of potential wells, suppressing the formation of small-scale structures and mitigating several associated problems [79, 450, 316, 449, 253]. A well-studied realization of warm DM is a thermal relic fermion that decoupled while still relativistic and later became non-relativistic as the universe expanded. Its free-streaming length, and thus the suppression of small-scale structure, depends primarily on the particle mass and decoupling temperature: a thermal relic with $m\sim(1\div 3)~\mathrm{keV}$ produces a cutoff in the matter power spectrum at comoving scales $\sim(0.1\div 1)~\mathrm{Mpc}$. Non-thermal scenarios—such as sterile neutrinos $\nu_{s}$ produced via the Dodelson–Widrow mechanism [161]—yield similar effects, though with distinct momentum distributions. Current bounds from the Lyman-$\alpha$ forest constrain the thermal relic mass to $m\gtrsim(3\div 5)~\mathrm{keV}$ [449, 253, 393, 204], while X-ray searches for the radiative decay $\nu_{s}\to\nu\gamma$ impose complementary bounds on sterile neutrino models [89]. Despite these stringent limits, warm DM remains an appealing minimal extension of CDM capable of mitigating small-scale structure tensions.
Another model motivated by these discrepancies is self-interacting DM [116, 424] in which, unlike warm DM, the astrophysical phenomenology is primarily controlled by the DM scattering cross section per unit mass $\sigma/m$, rather than the particle mass itself. Self-interactions redistribute DM from the inner to the outer regions of halos, softening the central cusp of the NFW profile and suppressing small-scale structure formation [377, 392, 261, 134, 386]. This model can reproduce the observed diversity in rotation curve shapes of dwarf galaxies [261, 134, 386] provided that the required amplitude for self-interactions is typically $\sigma/m\sim\mathcal{O}(0.1\div 10)\ \mathrm{cm^{2}/g}$ on dwarf galactic scales (with relative velocities $v\sim\mathcal{O}(10)\,\mathrm{km/s}$) whereas, to remain consistent with larger cluster-scale probes, the cross section has to be $\sigma/m\lesssim(0.1\div 1)\ \mathrm{cm^{2}/g}$ at $v\sim 10^{3}\ \mathrm{km/s}$, motivating velocity-dependent interactions [382, 390, 229]. Particle physics realizations that naturally produce this velocity dependence include Yukawa forces mediated by light scalar or vector (dark photon) fields [315, 264], giving rise to models spanning DM masses in the MeV to TeV range. Numerical simulations of self-interacting DM including baryonic feedback confirm that such models can mitigate small-scale tensions of the CDM paradigm [261, 134, 386, 264], while being compatible with existing astrophysical and cosmological constraints [439].
Further well-motivated class of DM candidates are axions and axion-like particles (ALPs), which arise as pseudo Nambu-Goldstone bosons associated with the spontaneous breaking of a global $U(1)$ symmetry [375, 454, 457]. The QCD axion was originally proposed to solve the strong CP problem, but it can also constitute all or part of the DM in the universe. ALPs generalize this idea to a broader class of light pseudoscalars, which need not solve the strong CP problem but can share similar properties, such as extremely weak couplings to Standard Model (SM) particles and long lifetimes [255, 329]. Both are typically very light, with masses ranging from $m_{a}\sim 10^{-12}\ \mathrm{eV}$ to $10^{-2}\ \mathrm{eV}$, depending on the symmetry-breaking scale $f_{a}$. In the simplest scenario, they are produced non-thermally via the vacuum misalignment mechanism [381, 2, 158], in which the initial field displacement from the minimum of its potential leads to coherent oscillations that behave as CDM. For QCD axions, the observed abundance is achieved for $f_{a}\sim(10^{11}\div 10^{12})\ \mathrm{GeV}$, corresponding to $m_{a}\sim 10^{-5}\ \mathrm{eV}$ [453, 85], whereas ALPs can span a broader range of masses and couplings, potentially allowing for heavier sub-eV candidates [250, 329]. These candidates are characterized by a rich phenomenology, which spans from ultralight quantum-suppressed structures on galactic scales to cold and coherent fields that behave like conventional CDM, potentially accessible through experimental, astrophysical, and cosmological observables [329, 250]. Further constraints on axion DM models will be discussed later in the Review.
@@
Yet another scenario that has gained significant attention in recent years is fuzzy DM, due to its rich phenomenology on small scales [246, 250]. In this model, DM consists of ultralight bosons with masses lighter than $10^{-20}\,{\rm eV}$. The extremely small mass invalidates a classical fluid description and allows DM to form a condensate on galactic scales through the balance between quantum pressure and gravity. This wave-like behavior gives rise to several distinctive phenomena, including the suppression of substructures below the de Broglie wavelength and the formation of cores in the inner regions of halos, with characteristic sizes set by quantum pressure and gravity [332, 333, 312]. On cosmological scales, however, fuzzy DM behaves similarly to CDM, preserving the large-scale successes of the standard model. Fuzzy DM can be naturally produced via non-thermal mechanisms in the early universe, giving rise to Bose-Einstein condensates with extremely long coherence lengths, leading to macroscopic wave interference effects in galactic halos. These quantum features lead to characteristic observational signatures, including granular density fluctuations, soliton-halo transitions, and modifications to the halo mass function at dwarf-galaxy scales. Observational constraints on fuzzy DM come from multiple probes. Lyman-$\alpha$ forest observations place a lower bound of $m>2\times 10^{-20}$ eV at 95% confidence level [254, 24, 281, 393]. Additional bounds arise from dwarf galaxy counts, stellar streams, and rotation curve measurements, which constrain the soliton core sizes and halo substructure [328, 398, 137, 40]. We defer the reader to recent reviews on fuzzy and Bose-Einstein condensate DM models for a discussion of further constraints on this scenario [427, 440, 194, 251, 173, 126].
@@
The extremely light mass range considered in fuzzy DM models requires that gravitational collapse be counterbalanced by quantum pressure. However, one can envision alternatives involving heavier DM candidates, extending up to the eV mass range. A prime example is the superfluid DM model, which postulates sub-eV DM particles with self-interactions capable of undergoing a superfluid phase transition in galactic halos. [^1] The formation of superfluids in galaxies can also be achieved through fermions with attractive interactions via Bardeen-Cooper-Schrieffer pairing. Such models typically involve a phase transition from a relativistic fermion fluid in the early Universe to a non-relativistic condensate at late times, resulting in a mildly evolving effective equation of state and characteristic astrophysical signatures, including core-like halo structures. The presence of a fermionic condensate can relax the Tremaine–Gunn lower bound [436] on the DM particle mass and can be observationally distinguished from ultralight bosonic DM through phenomena analogous to superconductivity, such as Andreev reflection and superconducting vortices. For developments in this direction, see Refs. [11, 12, 13, 115, 433, 202, 307, 9]. This framework allows for the formation of coherent superfluid cores, stabilized by the balance between self-interactions and gravity [215, 63, 64]. It offers a distinct approach to addressing the small-scale challenges mentioned earlier and admits, in principle, microscopic realizations. [^2] Another ultralight DM model, named self-interacting fuzzy DM, shares the same mass range as standard fuzzy DM but incorporates weak self-interactions [205, 345, 368, 111]. The main difference with SDM lies in the coherence length of the system. Although both frameworks involve self-interacting ultralight particles, self-interactions in the fuzzy case are not efficient enough to bring the system to equilibrium. As a result, the coherence length remains approximately set by quantum pressure, as in the non-interacting fuzzy DM scenario, in contrast to SDM, where self-interactions govern the macroscopic coherence properties. For earlier reviews on the SDM scenario, see Refs. [427, 194]. A schematic summary of the DM models discussed is shown in Fig. 1.
@@
In this review, we focus on the SDM scenario, summarizing its distinctive phenomenology on galactic scales and its potential to address small-scale challenges. Section 2 introduces the phenomenon of superfluidity, outlining its theoretical foundations within the frameworks of effective field theory and hydrodynamics. We also discuss finite-temperature effects, the interaction of the superfluid with external perturbations, and the role of gravity. Section 3 focuses on DM superfluidity in galactic environments, with particular attention to the formation of superfluid halos and their observable phenomenology. Section 4 explores its cosmological aspects, including possible production mechanisms and the subsequent evolution of the condensate. Section 5 provides an overview of the phenomenological implications of the SDM framework in both astrophysical and cosmological contexts, with particular emphasis on the dynamics of merging galaxy clusters, galaxy rotation curves and galactic relations, portals to the SM, vortex formation within halos and the development of DM spikes around BHs, while Section 6 examines the phenomenon of dynamical friction and its suppression in the presence of a superfluid, in contrast to the expectations from standard CDM. We present the theoretical foundations of this effect using both a quasiparticle description and a hydrodynamical approach. An application to the evolution of BH binaries is also discussed. Section 7 explores the possibility of long-range forces arising between baryons immersed in the DM superfluid, by coupling them to the constituent degrees of freedom. We present two illustrative examples capable of generating such interactions. Finally, Section 8 summarizes the main results of the review and provides a brief comparison between this model and other DM candidates.
@@ 2 Superfluidity and its effective description
where $c_{s}$ is the sound speed of the superfluid, expressed in terms of its pressure $P$ and density $\rho$. [^3] As usual, $\omega_{k}$ and $k$ denote the energy and momentum of the wave, respectively.
@@
where $E_{\text{\tiny kin}}$ is the kinetic energy of the superfluid. [^4] In the rest frame of the unperturbed fluid, the total energy is determined solely by the perturbation energy $\omega_{k}$. When transforming back to the laboratory frame via a Galilean boost, the energy acquires a contribution from the kinetic energy $E_{\text{\tiny kin}}$ of the moving fluid, one from the intrinsic energy $\omega_{k}$ of the perturbations in the comoving frame, and a Doppler term, $\vec{k}\!\cdot\!\vec{v}_{s}$, arising from boosting the perturbation wave packet to the laboratory frame. For the superfluid to be slowed down by the passage of the perturber, a certain amount of kinetic energy has to be damped by excited phonons. This is possible if the sum of the second and third terms of Eq. (2.2) is negative. By plugging in the phonon spectrum in Eq. (2.1), we have dissipation if $c_{s}k+\vec{k}\cdot\vec{v}_{s}\,<0$, which requires
@@ 2.1 Simplest example: superfluidity in a quartic theory
Here, $v$ describes the vacuum expectation value of the field, while $\mu$ is the relativistic chemical potential of the system, which is chosen in such a way that the field configuration (2.5) satisfies the equation of motion. In the non-relativistic limit, generically obtained by assuming $\lambda_{4}v^{2}\ll m^{2}$ [68], we can expand $\mu$ to obtain [^6] This is because, upon restoring the speed of light $c$, the chemical potential satisfies $\mu^{2}=m^{2}c^{4}+\lambda_{4}v^{2}c^{2}$. Expanding this expression in the non-relativistic (large-$c$) limit yields Eq. (2.6).
@@
Several approximations can be considered to simplify the analysis. First, since we aim to study non-relativistic superfluids, we can drop the $\dot{h}^{2}$ and $\dot{\pi}^{2}$ terms. [^7] This assumption is justified by the fact that those contributions are suppressed by an explicit factor of $c^{-2}$. The kinetic mixing term, however, is not suppressed, $2\mu c^{-2}\dot{\pi}h\sim 2m\dot{\pi}h$, and therefore survives in the non-relativistic limit. This implies that the typical energy scales of interest are smaller than the mass $m$ of the scalar. Second, we can consider a leading-order approximation in spatial derivatives of $h$ by dropping the gradient term $\big(\vec{\nabla}h\big)^{2}$. This is possible if we are interested in energy scales smaller than $\mu_{\text{\tiny NR}}$. In the context of DM superfluidity, the shortcomings of this approximation were discussed for dissipative processes [68] and in the analysis of the gravitational stability of the superfluid phase [57].
@@
Lastly, we comment on the requirement of repulsive self-interactions for superfluidity. If the theory had negative quartic coupling, $\lambda_{4}<0$, the long-wavelength fluctuations of the homogeneous solution would have an imaginary sound speed (see Eq. (2.11)) and thus would be unstable. One may therefore consider the simple quartic potential with repulsive self-interactions as in (2.4). However, such realizations are typically sensitive to sizable radiative corrections, which can destabilize the tiny mass required for superfluidity on astrophysical scales.
This radiative instability can be avoided if the potential is protected by an almost exact symmetry. A well-known way to stabilize it is to endow the scalar field with a shift symmetry. This naturally arises if the scalar is the Nambu-Goldstone boson associated with the spontaneous breaking of a global symmetry. The shift symmetry restricts the scalar to appear only through derivative interactions. Consequently, a small explicit breaking of the global symmetry is required to introduce non-derivative terms, generating a mass and self-interactions. In this case, all such couplings are proportional to the parameter controlling the explicit symmetry breaking, the spurion, which also ensures the radiative stability of the potential as in (2.4).
However, generating a repulsive potential through this mechanism is non-trivial. For instance, in the well-known case of the QCD axion, where the shift symmetry is broken by instanton effects, the resulting potential is a periodic cosine. Expanding it in powers of the scalar field, the quartic term carries an attractive coupling [441] (see also Sec. 5.4 for further details on the QCD axion):
$V(a)=f_{a}^{2}m_{a}^{2}\left[1-\cos\left(\frac{a}{f_{a}}\right)\right]\sim\frac{m_{a}^{2}}{2}a^{2}-\frac{m_{a}^{2}}{4!f_{a}^{2}}a^{4}+\ldots$ (2.12)
Here, $f_{a}$ is the decay constant of the axion, while $m_{a}$ is the spurion. As a result, such particles cannot produce long-range correlations due to their attractive self-interactions [225]. More generally, pseudo-Nambu-Goldstone bosons arising from the breaking of a compact group typically have periodic trigonometric potentials, and any single cosine or sine term inevitably leads to attractive interactions.
The challenges of constructing pseudo-Nambu-Goldstone boson models with repulsive self-interactions are discussed in [189], where various scenarios are analyzed. The author shows that, in most cases, repulsive self-interactions cannot be achieved. On the other hand, they also show a working example in which the repulsive interactions arise from an axion originating as the remnant of a five-dimensional $U(1)$ gauge symmetry.
These limiting considerations, however, arise from naturalness arguments for the model parameters, analogous to the so-called hierarchy problem for the Higgs boson mass. Although this is an interesting issue worthy of further investigation, at the level of theoretical consistency the construction of repulsively self-interacting massive bosons requires no more effort than the Higgs mechanism — the key difference being that no tachyonic mass term is needed.
@@ 2.4 Gross–Pitaevskii equation and the hydrodynamical formulation
Substituting back into (2.31), and ignoring quadratic in time derivatives, we obtain the Schrödinger-like equation,
@@
Before proceeding, let us mention that the derivation of Eq. (2.33) relied on further few assumptions outlined in Refs. [399, 400]. In particular, the decomposition (2.32) allows to rewrite the relativistic field $\Phi$ in terms of its nonrelativistic component $\Psi$, which is a slowly varying function of time and space (compared to the dominant frequency of the system given by the scalar field mass $m$). Due to the quartic nonlinearities involved in the relativistic system (2.31), small amplitude, high-frequency oscillations sourced nonlinearly by $\Psi$ would in turn backreact on $\Psi$ itself, affecting its dynamics. This effect can be taken care of by integrating out the fast modes, thus inducing relativistic corrections to the Gross–Pitaevskii equation, which we neglect in this section and postpone to Sec. 2.7. Furthermore, in an expanding background, there would be additional terms in Eq. (2.33) proportional to both the metric and matter perturbation sectors, which we also neglect as small corrections in the late universe, when the fields are essentially nonrelativistic.